# Noemi Walczak Université Paris 1 Panthéon-Sorbonne, Paris, France Kenan Huremović Armando Rungi

### v3

@@ A.2 The model
 $\qquad\text{ s.t. }q_{i}^{j}\leq\Upsilon_{i}^{j}\left[A_{i}^{j}(l_{i}^{j})^{\beta_{i}^{j}}\left(\sum_{n\in\mathcal{N}}\sum_{k\in\mathcal{J}}(\iota_{ni}^{kj})^{\frac{1}{\theta}}(z_{ni}^{kj})^{\frac{\theta-1}{\theta}}\right)^{\frac{\theta}{\theta-1}(1-\beta_{i}^{j})}\right]^{1-\rho_{i}^{j}}[e_{i}^{j}]^{\rho_{i}^{j}}$
@@
$\begin{cases}\frac{\partial\mathcal{L}}{\partial l_{i}^{j}}=w_{i}-\lambda(1-\rho_{i}^{j})\beta_{i}^{j}\frac{q_{i}^{j}}{l_{i}^{j}}=0\\ \frac{\partial\mathcal{L}}{\partial e_{i}^{j}}=t_{i}(1-\epsilon_{i}^{j})-\lambda\rho_{i}^{j}\frac{q_{i}^{j}}{e_{i}^{j}}=0\\ \forall n,k:\quad\frac{\partial\mathcal{L}}{\partial z_{ni}^{kj}}=\frac{\partial\mathcal{L}}{\partial M_{i}^{j}}\frac{\partial M_{i}^{j}}{\partial z_{ni}^{kj}}=p_{ni}^{kj}-\lambda\left[(1-\beta_{i}^{j})(1-\rho_{i}^{j})\frac{q_{i}^{j}}{M_{i}^{j}}\right]\left[(\iota_{ni}^{kj})^{\frac{1}{\theta}}(z_{ni}^{kj})^{\frac{\theta-1}{\theta}-1}(M_{i}^{j})^{\frac{1}{\theta}}\right]=0\\ \frac{\partial\mathcal{L}}{\partial\lambda}=q_{i}^{j}-\Upsilon_{i}^{j}\left[A_{i}^{j}(l_{i}^{j})^{\beta_{i}^{j}}(M_{i}^{j})^{1-\beta_{i}^{j}}\right]^{1-\rho_{i}^{j}}[e_{i}^{j}]^{\rho_{i}^{j}}=0\end{cases}$
@@
We define the steady state as the combination of equilibrium prices and quantities in an undistorted economy with no wedges, where $(1-\epsilon_{n}^{k})^{\rho_{n}^{k}}(A_{n}^{k})^{-(1-\rho_{n}^{k})}=1\quad\forall i,n\in\mathcal{N}\ j,k\in\mathcal{J}$. From Lemma A.2. and the pricing rule we have:
@@
And the system of equations is solved for $p_{i}^{j}=w_{i}=t_{i}=1\quad\forall i,j$.
@@
Proof. Emissions embodied in trade are calculated using the vector of production-based emissions and input-output multipliers. We borrow the definition from the OECD [30] and we adapt it to align with the structure of our model. The emissions embodied in imports are given by the following vector $\mathbf{EEI}\in\mathbb{R}^{NJ}$ with entries corresponding to the emissions generated from the production of each good $j\in\mathcal{J}$ in every country $i\in\mathcal{N}$ embodied in imports from all trade partners. In particular, the first equivalence in equation (17) corresponds to the OECD definition, while the subsequent equalities restates it in terms of our notation and modeling framework:
@@ A.3 Additional Figures
Figure A3: 2018 Share of EU Carbon Intensive Imports by country in Total Intermediate Purchases. Highlighted, the 10 biggest exporters of carbon intensive goods to the EU
@@
Figure A6: Policy-induced changes in European Sales Shares of selected sectoral goods (% of baseline level) – $\theta=8$
@@ A.4 Additional Tables
Table A2: Descriptive statistics by country of carbon intensities ($\rho_{i}^{j}$) in carbon-intensive sectors.
Variable Current CBAM Full CBAM
$\theta=2$ $\theta=8$ $\theta=2$ $\theta=8$
Panel (a) Average $\Delta\%$ in EU Share of Foreign Purchases
Total
-0.20
-1.25
-0.37
-2.88
   …of clean intermediate goods
0.23
1.06
0.82
3.67
   …of dirty intermediate goods
-0.37
-2.17
-0.70
-4.66
Panel (b) Average $\Delta\%$ in EU Share of Domestic Purchases
@@
   …of dirty intermediate goods 0 0.07 0.01 0.20
Panel (c) Average $\Delta\%$ in EU Domar Weights
Total
0.09
-0.05
0.34
-0.06
   …of clean intermediate goods
0.14
0.08
0.45
0.15
   …of dirty intermediate goods
0.06
-0.12
0.27
-0.18
Table A3: Policy-induced changes in EU-ETS shares (% of baseline) for $\theta=2$ and $\theta=8$. Values are country means; standard deviations in parentheses where applicable.
Variable Current CBAM Full CBAM
@@
Table A4: Policy-induced changes in Emissions Embodied in Imports (% of baseline level) for $\theta=2$ and $\theta=8$. Values are country means; standard deviations in parentheses where applicable.

### v4

@@ A.2 The model
 $\hskip 18.49988pt\text{ s.t. }q_{i}^{j}\leq\Upsilon_{i}^{j}\left[A_{i}^{j}(l_{i}^{j})^{\beta_{i}^{j}}\left(\sum_{n\in\mathcal{N}}\sum_{k\in\mathcal{J}}(\iota_{ni}^{kj})^{\frac{1}{\theta}}(z_{ni}^{kj})^{\frac{\theta-1}{\theta}}\right)^{\frac{\theta}{\theta-1}(1-\beta_{i}^{j})}\right]^{1-\rho_{i}^{j}}[e_{i}^{j}]^{\rho_{i}^{j}}$
@@
$\begin{cases}\frac{\partial\mathcal{L}}{\partial l_{i}^{j}}=w_{i}-\lambda(1-\rho_{i}^{j})\beta_{i}^{j}\frac{q_{i}^{j}}{l_{i}^{j}}=0\\ \frac{\partial\mathcal{L}}{\partial e_{i}^{j}}=t_{i}(1-\epsilon_{i}^{j})-\lambda\rho_{i}^{j}\frac{q_{i}^{j}}{e_{i}^{j}}=0\\ \forall n,k:\qquad\frac{\partial\mathcal{L}}{\partial z_{ni}^{kj}}=\frac{\partial\mathcal{L}}{\partial M_{i}^{j}}\frac{\partial M_{i}^{j}}{\partial z_{ni}^{kj}}=p_{ni}^{kj}-\lambda\left[(1-\beta_{i}^{j})(1-\rho_{i}^{j})\frac{q_{i}^{j}}{M_{i}^{j}}\right]\left[(\iota_{ni}^{kj})^{\frac{1}{\theta}}(z_{ni}^{kj})^{\frac{\theta-1}{\theta}-1}(M_{i}^{j})^{\frac{1}{\theta}}\right]=0\\ \frac{\partial\mathcal{L}}{\partial\lambda}=q_{i}^{j}-\Upsilon_{i}^{j}\left[A_{i}^{j}(l_{i}^{j})^{\beta_{i}^{j}}(M_{i}^{j})^{1-\beta_{i}^{j}}\right]^{1-\rho_{i}^{j}}[e_{i}^{j}]^{\rho_{i}^{j}}=0\end{cases}$
@@
We define the steady state as the combination of equilibrium prices and quantities in an undistorted economy with no wedges, where $(1-\epsilon_{n}^{k})^{\rho_{n}^{k}}(A_{n}^{k})^{-(1-\rho_{n}^{k})}=1\qquad\forall i,n\in\mathcal{N}\ j,k\in\mathcal{J}$. From Lemma A.2. and the pricing rule we have:
@@
And the system of equations is solved for $p_{i}^{j}=w_{i}=t_{i}=1\qquad\forall i,j$.
@@
Proof. Emissions embodied in trade are calculated using the vector of production-based emissions and input-output multipliers. We borrow the definition from the OECD [yamano2020co2] and we adapt it to align with the structure of our model. The emissions embodied in imports are given by the following vector $\mathbf{EEI}\in\mathbb{R}^{NJ}$ with entries corresponding to the emissions generated from the production of each good $j\in\mathcal{J}$ in every country $i\in\mathcal{N}$ embodied in imports from all trade partners. In particular, the first equivalence in equation (17) corresponds to the OECD definition, while the subsequent equalities restates it in terms of our notation and modeling framework:
@@ A.3 Additional Figures
Figure A3: 2018 Share of EU Carbon Intensive Imports by country in Total Intermediate Purchases. Highlighted, the 10 biggest exporters of carbon-intensive goods to the EU
@@
Figure A6: Changes in EU sales shares of selected sectoral goods (% of baseline year 2024) – $\theta=8$
@@ A.4 Additional Tables
Table A2: Country-level descriptive statistics of carbon intensities ($\rho_{i}^{j}$) in carbon-intensive sectors.
Variable Reduced CBAM Full CBAM
Endog. Exog. Endog. Exog.
Mean St.dev. Mean St.dev. Mean St.dev. Mean St.dev.
Panel (a) $\Delta\%$ in EU Share of Foreign Purchases
Total -0.20 0.48 -1.25 2.48 -0.37 0.85 -2.88 5.99
    …of clean intermediate goods 0.23 0.14 1.06 0.61 0.82 0.81 3.67 0.61
    …of dirty intermediate goods -0.37 0.89 -2.17 4.48 -0.70 1.19 -4.66 7.44
Panel (b) $\Delta\%$ in EU Share of Domestic Purchases
@@
    …of dirty intermediate goods 0.00 0.07 0.01 0.20
Panel (c) $\Delta\%$ in EU Domar Weights
Total 0.09 0.11 -0.05 0.34 0.34 0.22 -0.06 0.71
    …of clean intermediate goods 0.14 0.06 0.08 0.14 0.45 0.18 0.15 0.47
    …of dirty intermediate goods 0.06 0.12 -0.12 0.39 0.27 0.22 -0.18 0.80
Table A3: Changes in EU-ETS shares (% baseline level year 2024) for changing elasticities, $\theta=2$ and $\theta=8$. Values are country means; standard deviations in parentheses where applicable.
Variable Reduced CBAM Full CBAM
@@
Table A4: Changes in emissions embodied in imports (% of baseline level year 2024) for $\theta=2$ and $\theta=8$. Values are country means; standard deviations in parentheses where applicable.